OpenStudy (anonymous):

Creative Landscaping has 60 yd of fencing with which to enclose a rectangular flower garden. If the garden is x yards long, express the garden’s area as a function of the length. I just can't figure out how to represent the width?

OpenStudy (anonymous):

the perimeter is 60 so, 60=2x+2w where w is the width. so, w=30-x and, area=x(30-x) or, A=30x-x^2

OpenStudy (anonymous):

Thank you so much! I knew what I was trying to say, but I couldn't seem to get it in math terms! :))

OpenStudy (anonymous):

np :)